A company that supplies ready-mix concrete receives, on average, six orders per day.



(a) What is the probability that, on a given day:


(i) only one order will be received?


(ii) no more than three orders will be received?


(iii) at least three orders will be received?


(b) What is the probability that, on a given half-day, only one order will be received?


(c) What is the mean and standard deviation of orders received per day?



1
Expert's answer
2022-06-16T14:32:09-0400

Solution:

Let count all possible outcomes and ways:.

Number of possible outcomes: "N=2^{n};" here "n" - number of orders.

Number of ways: "^{n}C_{m}=\frac{n!}{(n-m)!m!};" here "m" - number of received orders. So,

"^{6}C_{0}=\frac{6!}{(6-0)!0!}=1;"

"^{6}C_{1}=\frac{6!}{(6-1)!1!}=6;"

"^{6}C_{2}=\frac{6!}{(6-2)!2!}=15;"

"^{6}C_{3}=\frac{6!}{(6-3)!3!}=20;"

"^{6}C_{4}=\frac{6!}{(6-4)!4!}=15;"

"^{6}C_{5}=\frac{6!}{(6-5)!5!}=6;"

"^{6}C_{6}=\frac{6!}{(6-6)!6!}=1;"

"N=2^{6}=64;"

a)

1) "P(1)=\frac{6}{64}=0.09375;"

2) No more than three, so, we need to add terms with m from 0 to 3:

"P(m\le3)=P(0)+P(1)+P(2)+P(3)=\frac{1+6+15+20}{64}=0.65625;"

3) At least three order received, so, we need to add terms with m from 3 to 6;

"P(m\ge3)=P(3)+P(4)+P(5)+P(6)=\frac{20+15+6+1}{64}=0.65625;"

b) If a half day, it means "n=\frac{6}{2}=3;"

"N=2^{3}=8-" numbers of possible outcomes;

Number of ways:

"^{3}C_{0}=\frac{3!}{(3-0)!0!}=1;"

"^{3}C_{1}=\frac{3!}{(3-1)!1!}=3;"

"^{3}C_{2}=\frac{3!}{(3-2)!2!}=3;"

"^{3}C_{3}=\frac{3!}{(3-3)!3!}=1;"

"P(1)=\frac{3}{8}=0.375;"

c) In the binomial distribution:

"\mu=nP(1)=64\times0.09375=6" is the mean of the distribution;

"\sigma=\sqrt{nP(1)(1-P(1))}=2.33" is the standard deviation of the distribution.

Answer:

a)

1) "P(1)=0.09375;"

2) "P(m\le3)=0.65625;"

3) "P(m\ge3)=0.65625;"

b) "P(1)=0.375;"

c) "\mu=6;" "\sigma=2.33."